$79 million and issued a nine-month promissory note. Interest was discounted at issuance at On October 1, Eder Fabrication borrowed a 11% discount rate. Prepare the journal entry for the issuance of the note and the appropriate reporting period. (If no entry is required for a transaction/event, select your answers in whole dollars.) adjusting entry for the note at December 31, the end of the No journal entry required in the first account field. Enter View transaction list Journal entry worksheet 2 Record the issuance of the note. Note: Enter debits before credits. General Journal Debit Credit Event
Answer #1

Journal entry for issuance of note:

Cash

$ 724,82,500

Interest Expense

$ 65,17,500

       Notes Payable

$ 790,00,000

Discount= 790,00,000* 0.11*9/12= $ 65,17,500

Proceeds=790,00,000-65,17,500=$ 724,82,500

Recording Entry on December 31

Interest Expense

$ 2172500

       Interest Payable

$ 2172500